OpenStudy (anonymous):

Assume f(x) = -2x + 8 and g(x) = 5x. What is the value of (f - g)(2)?

OpenStudy (accidentalaichan):

We substitute g for x. f(5) = -2(5) + 8 Now we solve. f(5) = -10 +8 f= -2 So now we solve the second equation. (f-g)(2) (-2 - 5)(2)

OpenStudy (anonymous):

so ur answer is -14?

OpenStudy (accidentalaichan):

yes, That is what I have.